OpenStudy (anonymous):

Please help me! For each set of conditions, write an appropriate quadratic function rule: 1. X-intercepts at (-4,0) and (12,0) and opening downward 2. X- intercepts at (-2,0) and (6,0) with y intercept at (0,-60) 3. X-intercepts at (-1,0) and (3,0) with a minimum at (1,-4) Please help me!

OpenStudy (anonymous):

for the first one you can start with \[y=(x+4)(x-12)\] but since it opens down, you should make it \[y=-(x+4)(x-12)\]

OpenStudy (anonymous):

then i guess you have to multiply it out

OpenStudy (anonymous):

thank you!!! do you have any idea on how to do the other two?

OpenStudy (agent47):

x intercepts at (-2,0) and (6,0) mean: y = 0 when x=-2 and x=6, or: (x - (-2))*(x-6) = 0, or (x+2)(x-6)=0

OpenStudy (anonymous):

thnx :)

OpenStudy (agent47):

Now we just need the leading coefficient in front of those parentheses.. y = A(x+2)(x-6) when x=0, y=-60: -60 = A (-12), A=5.

OpenStudy (agent47):

So: y = 5(x+2)(x-6) <-simplify.

OpenStudy (anonymous):

ur first problem in a year